The following condensed income statements of the Jackson Holding Company are presented for the two years ended December 31, 2divisions assets on December 31 was $3,940,000. Prepare revised income statements according to generally accepted accountingdivisions assets on December 31 was $3,940,000. Prepare revised income statements according to generally accepted accountingdivisions assets on December 31 was $3,940,000. Prepare revised income statements according to generally accepted accounting
Answer #1

Dear Student,

If any doubts then feel free to ask.

If satisfied with the answer, then kindly rate the answer.

Part 1

JACKSON HOLDING COMPANY

Comparative Income Statements (in part)

For the Years Ended December 31

2018

2017

Income from continuing operations before income taxes

2220000

720000

Income tax benefit (expense)

(888000)

(288000)

Income from continuing operations

1332000

432000

Discontinued operations gain (loss):

Income (loss) from operations of discontinued component

1060000

320000

Income tax benefit (expense)

(42400)

(128000)

Income (loss) on discontinued operations

963750

228750

Net income

2349600

624000

2220000*40% =888000

720000*40% = 288000

1060000*40% = 424000

320000*40% = 128000

2018

2017

Income from continuing operations before income taxes:

Unadjusted

2640000

1040000

Less : Income from discontinued operations

420000

320000

Adjusted

2220000

720000

2018

2017

Income from discontinued operations:

Income from operations

420000

320000

Gain on disposal

640000

Total

1060000

320000

Part 2

JACKSON HOLDING COMPANY

Comparative Income Statements (in part)

For the Years Ended December 31

2018

2017

Income from continuing operations before income taxes

2220000

720000

Income tax benefit (expense)

(888000)

(288000)

Income from continuing operations

1332000

432000

Discontinued operations gain (loss):

Income (loss) from operations of discontinued component

420000

320000

Income tax benefit (expense)

(168000)

(128000)

Income (loss) on discontinued operations

252000

228750

Net income

1584000

624000

420000*40% = 168000

Part 3

JACKSON HOLDING COMPANY

Comparative Income Statements (in part)

For the Years Ended December 31

2018

2017

Income from continuing operations before income taxes

2220000

720000

Income tax benefit (expense)

(888000)

(288000)

Income from continuing operations

1332000

432000

Discontinued operations gain (loss):

Income (loss) from operations of discontinued component

(120000)

320000

Income tax benefit (expense)

48000

(128000)

Income (loss) on discontinued operations

(72000)

228750

Net income

1260000

624000

420000-(4480000-3940000)= (120000)

120000*40% = 48000